OpenStudy (anonymous):

Help please to make sure I am correct :) A circle that originally had a circumference of 31.4 cm, now has a circumference of 15.7 cm. What happened to the radius?

OpenStudy (anonymous):

Is the radius halved?

OpenStudy (anonymous):

Yep your corect! ;) Nice job~

OpenStudy (cwrw238):

the relation between the circumference and the radius is C = 2 * pi * r where pi is a constant

OpenStudy (anonymous):

Yea, thank u! :)

OpenStudy (anonymous):

makes sense some how lol

OpenStudy (cwrw238):

31.4 = 2pi r making r = 31.4 / 2 pi 15.7 = 2 pi r making r = 15.7 / 2 pi which is a half of 31.4 / 2 pi